### v1

# On the Exteneded Painlevé Property of Riccati Hiearchies

Abstract
This article presents a new expression for Riccati chains using Bell polynomials. The quasi-homogeneous properties of these chains are explored, and the Kovalevskaya exponents are analyzed. Additionally, we demonstrate that Riccati chains exhibit the extended Painlevé property, employing the method of weighted projective space.
1 Introduction
The Riccati equation is a fundamental nonlinear differential equation that serves as the simplest nonlinear analog of the second-order linear equation. Its higher-order generalizations, known as Riccati chains, arise naturally when taking successive covariant derivatives with respect to the logarithmic derivative $v=\frac{f_{x}}{f}$ of a projective field $f$. These chains form a hierarchy that links linear, nonlinear, and integrable systems through the operator $L=\frac{\partial}{\partial x}+v$ [1, 2].
Riccati chains exhibit properties crucial for integrability, such as quasi-homogeneity. A vector field $f=(f_{1},\dots,f_{m})$ is quasi-homogeneous if it satisfies a scaling property under the action of a C*-group, a feature that influences both the dynamics and integrability of the system. This quasi-homogeneous structure is important in the study of the Painlevé property, which ensures that the solutions of the equations do not have movable singularities [3, 8]. In this work, we investigate the extended Painlevé property for Riccati chains using the method of weighted projective space, which is a powerful tool for studying the integrability of differential systems [5, 6, 7].
To facilitate the study of these chains, we express the Riccati chain equations using Bell polynomials, which provide a compact and efficient representation. Bell polynomials capture the recursive structure of Riccati chains, making the analysis of their singularities more tractable. Additionally, we compute the Kovalevskaya exponents, which are essential for classifying the singularities of these systems and determining their integrability.
In this article, we demonstrate that Riccati chains satisfy the extended Painlevé property, showing that they are integrable in the extended sense. This result builds upon earlier work on the Painlevé equations and the classification of integrable systems [12, 13].
The article is structured as follows: Section 2 introduces the theoretical framework and formulation of the Riccati chain equations. Section 3 discusses the quasi-homogeneity and weight classification of Riccati chains. Section 4 presents the computation of Kovalevskaya exponents and the analysis of indicial loci. Section 5 proves that Riccati chains possess the extended Painlevé property using the weighted projective space method. We conclude in Section 6 with a discussion of the implications of our results and potential future directions.
2 Theoretical Setting
In this section, we present the theoretical framework for the Riccati chain, focusing on its recursive structure and the use of Bell polynomials for a compact representation.
Definition 2.1.
Let $L$ be the first-order differential operator
$L=\frac{d}{dz}+cu(z),\quad c\in\mathbb{R}^{*},$
where $u(z)$ is a potential function. The $m$-order Riccati chain equation is a generalization of the standard Riccati equation and can be written as:
$L^{m}u(z)+\sum_{j=1}^{n}\alpha_{j}(z)L^{j-1}u(z)+\alpha_{0}(z)=0,$
where $m\in\mathbb{N}$, and the coefficient functions $\alpha_{j}(z)$ may depend on $u(z)$ and its derivatives.
Remark 2.2.
The operator $L^{m}u$ expands recursively as
$L^{k}u=\left(\frac{d}{dz}+u\right)L^{k-1}u,$
generating a polynomial in $u$ and its derivatives.
The recursive nature of this system suggests a way to express the Riccati chain efficiently using Bell polynomials, which offer a systematic way to represent the terms in the Riccati chain. The following definitions and propositions provide the necessary tools for this representation.
Definition 2.3 (Exponential partial Bell polynomial).
For integers $m\geq k\geq 1$, the exponential partial Bell polynomial $B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})$ is defined by the generating function
$\frac{1}{k!}\!\left(\sum_{n=1}^{\infty}x_{n}\,\frac{t^{n}}{n!}\right)^{\!k}=\sum_{m=k}^{\infty}B_{m,k}(x_{1},x_{2},\dots,x_{m-k+1})\,\frac{t^{m}}{m!}.$
Equivalently, let
$\mathcal{J}_{m,k}=\Bigl\{(j_{1},\dots,j_{m-k+1})\in\mathbb{N}_{0}^{\,m-k+1}\;\Big|\;\begin{array}[]{l}j_{1}+\cdots+j_{m-k+1}=k,\\[3.0pt] 1j_{1}+2j_{2}+\cdots+(m-k+1)j_{m-k+1}=m\end{array}\Bigr\}.$
Then
$B_{m,k}(x_{1},\dots,x_{m-k+1})=\sum_{(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}}\frac{m!}{j_{1}!\,j_{2}!\cdots j_{m-k+1}!}\left(\frac{x_{1}}{1!}\right)^{\!j_{1}}\left(\frac{x_{2}}{2!}\right)^{\!j_{2}}\!\cdots\left(\frac{x_{m-k+1}}{(m-k+1)!}\right)^{\!j_{m-k+1}}.$
Each $(j_{1},\dots,j_{m-k+1})\in\mathcal{J}_{m,k}$ corresponds to a way of partitioning $m$ into $k$ parts, where $j_{i}$ denotes the number of parts of size $i$.
Example 2.4.
Some examples of the Bell polynomials are:
$B_{1,1}$ $=x_{1},$
$B_{2,1}$ $=x_{2},\quad B_{2,2}=x_{1}^{2},$
$B_{3,1}$ $=x_{3},\quad B_{3,2}=3x_{1}x_{2},\quad B_{3,3}=x_{1}^{3},$
$B_{4,1}$ $=x_{4},\quad B_{4,2}=4x_{1}x_{3}+3x_{2}^{2},\quad B_{4,3}=6x_{1}^{2}x_{2},\quad B_{4,4}=x_{1}^{4}.$
These polynomials provide the building blocks for the Bell-polynomial expansion of the Riccati chain.
Proposition 2.5 (Bell–polynomial form of the Riccati chain).
Fix $c\in\mathbb{R}$ and write $L:=\partial_{z}+c\,u$. For $m\in\mathbb{N}$ and smooth coefficients $\alpha_{j}(z)$, the $m$-th Riccati–chain equation
$L^{\,m}u(z)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\,L^{\,j-1}u(z)\;+\;\alpha_{0}(z)\;=\;0$
is equivalent to
$\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big)\;+\;\sum_{j=1}^{m}\alpha_{j}(z)\!\sum_{k=1}^{j}c^{\,k-1}\,B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big)\;+\;\alpha_{0}(z)\;=\;0,$
where $B_{n,k}$ are the exponential partial Bell polynomials.
Proof.
The standard identity for iterates of $L=\partial_{z}+c\,u$ gives
$L^{\,m}u=\sum_{k=1}^{m+1}c^{\,k-1}\,B_{m+1,k}\!\big(u,u^{\prime},\dots,u^{(m)}\big),$
which follows by induction from the recursion $L^{\,m+1}u=(\partial_{z}+cu)\,L^{\,m}u$ and the Bell–polynomial recurrence
$B_{n+1,k}(_{1},\dots,x_{n+1})=\sum_{r=0}^{n-k+1}\binom{n}{r}\,x_{r+1}\,B_{n-r,k-1}(x_{1},\dots,x_{n-r}).$
Applying the same expansion to each $L^{\,j-1}u$ yields
$L^{\,j-1}u=\sum_{k=1}^{j}c^{\,k-1}B_{j,k}\!\big(u,u^{\prime},\dots,u^{(j-1)}\big).$
Substituting these two identities into the Riccati–chain equation gives the stated Bell–polynomial form. ∎
Example 2.6 (Case $m=3$).
With $L=\partial_{z}+c\,u$, the Bell–polynomial expansion gives
$L^{3}u=\sum_{k=1}^{4}c^{k-1}B_{4,k}(u,u^{\prime},u^{\prime\prime},u^{\prime\prime\prime})=u^{\prime\prime\prime}+4c\,u\,u^{\prime\prime}+3c\,(u^{\prime})^{2}+6c^{2}u^{2}u^{\prime}+c^{3}u^{4},$
since
$B_{4,1}=u^{\prime\prime\prime},\quad B_{4,2}=4u\,u^{\prime\prime}+3(u^{\prime})^{2},\quad B_{4,3}=6u^{2}u^{\prime},\quad B_{4,4}=u^{4}.$
Similarly,
$L^{2}u=\sum_{k=1}^{3}c^{k-1}B_{3,k}(u,u^{\prime},u^{\prime\prime})=u^{\prime\prime}+3c\,u\,u^{\prime}+c^{2}u^{3},\qquad Lu=u^{\prime}+c\,u^{2}.$
Substituting into
$L^{3}u+\alpha_{3}(z)\,L^{2}u+\alpha_{2}(z)\,Lu+\alpha_{1}(z)\,u+\alpha_{0}(z)=0$
yields
$u^{\prime\prime\prime}$ $+\big(\alpha_{3}(z)+4c\,u\big)u^{\prime\prime}+3c\,(u^{\prime})^{2}+\big(6c^{2}u^{2}+3c\,\alpha_{3}(z)\,u+\alpha_{2}(z)\big)u^{\prime}$
$\qquad+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\,\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)\;=0,$
which matches the $m=3$ member of the Riccati chain.
3 Quasi-homogeneity and Weights
Let us consider the system of differential equations
$\frac{dx_{i}}{dz}=f_{i}(x_{1},\cdots,x_{m},z),\quad i=1,\cdots,m,$ (3.1)
where the functions $f_{i}$ are polynomials in $(x_{1},\cdots,x_{m},z)\in\mathbb{C}^{m+1}$.
The key feature of these equations that we explore is quasi-homogeneity, a property that influences the structure and behavior of the solutions. A vector field $(f_{1},\cdots,f_{m})$ is said to be quasi-homogeneous if it satisfies the following scaling property:
$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{p_{i}+1}f_{i}(x_{1},\dots,x_{m},z),$
for any $\lambda\in\mathbb{C}$ and $i=1,\dots,m$, where $(p_{1},\cdots,p_{m},r)\in\mathbb{Z}^{m+1}_{>0}$ is a set of positive weights. The integers $p_{i}$ and $r$ are referred to as the weighted degrees of $x_{i}$ and $z$, respectively. This scaling property is crucial because it governs the dynamics of the system and helps classify its solutions.
The Newton polyhedron of the system (3.1) is defined as the convex hull of the union of the positive orthants in $\mathbb{R}^{m+1}$, with vertices at the exponents of the monomials in the system. The Newton diagram, which is the union of the compact faces of the Newton polyhedron, provides valuable information about the system’s solutions and their singularity structure.
If the Newton diagram consists of a single compact face, we can introduce a tuple of integers $(p_{1},\dots,p_{m},r,s)$ such that the compact face lies on the plane
$p_{1}x_{1}+\cdots+p_{m}x_{m}+rz=s$
in $\mathbb{R}^{m+1}$. In this case, the function $f_{i}$ satisfies the scaling relation:
$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{s-r+p_{i}}f_{i}(x_{1},\dots,x_{m},z),$
for any $\lambda\in\mathbb{C}$.
Next, we introduce the $m$-th Riccati chain equation. By the map $x_{k}\rightarrow L^{k-1}u$, the Riccati chain can be written as a system of first-order differential equations:
$x_{k}^{\prime}$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (3.2)
$x_{m}^{\prime}$ $=-\sum_{j=0}^{m}\alpha_{j}(z)x_{j}-cx_{1}x_{m},\quad$ $(x_{0}\equiv 1),$
which reduces to the standard Riccati equation when $m=1$:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
The structure of the Riccati chain can be better understood by considering the weights of its terms. In particular, the weights of the terms $\alpha_{j}(z)$ are related to the scaling properties of the system. For example, the terms $\alpha_{j}(z)$ associated with the Riccati chain can be written as powers of $z$ with the following weights:
[…]
These weights play an essential role in understanding the quasi-homogeneous structure of the Riccati chain.
Weight Classification for the Riccati Chain: The weights of the terms $\alpha_{j}(z)$ are crucial for classifying the Riccati chain’s solutions. Specifically, for the $m$-th Riccati chain given by (3.2), the weights are given by:
$p_{k}=k\quad(1\leq k\leq m),\quad r\in\{1,\cdots,m+1\},\quad s=r+1.$
Assuming the full vector field $F=(f_{1},\dots,f_{m})$ is quasi-homogeneous of degree +1+1, the coefficient functions $\alpha_{j}(z)$ satisfy the following weight classification:
$\text{wt}(\alpha_{j})=m+1-j,\quad 0\leq j\leq m.$
In particular, if $\alpha_{j}(z)$ are polynomials in $z$, they take the form
$\alpha_{j}(z)=\begin{cases}A_{j}z^{\frac{m-j+1}{r}}&\text{if}\quad\frac{m-j+1}{r}\in\mathbb{Z},\\ 0&\text{otherwise}.\end{cases}$
This weight classification helps in understanding the structure of the Riccati chain and its integrability.
Examples for $n=1$
For $n=1$, the Riccati chain reduces to the standard Riccati equation:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
The weight classification of this equation gives two types of quasi-homogeneous solutions:
$x_{1}^{\prime}=-z+cx_{1}^{2}\quad\text{(Airy type)},$
and
$x_{1}^{\prime}=-zx_{1}-z^{2}+cx_{1}^{2}\quad\text{(Hermite type)}.$
The Airy-type solution is discussed in [4], where a compactified Riccati equation of Airy type on a weighted projective space is studied.
Higher-order Examples and Chazy–Bureau Equations
For higher-order cases, the Riccati chain leads to more complex equations. For $n=3$, the Riccati chain in scalar form reads:
$u^{\prime\prime\prime}$ $+(\alpha_{3}(z)+4cu)u^{\prime\prime}+3c(u^{\prime})^{2}$
$+\left(6c^{2}u^{2}+3c\alpha_{3}(z)u+\alpha_{2}(z)\right)u^{\prime}$
$+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)=0.$
For $\alpha_{j}(z)=0$ and $c=1$, this equation simplifies to the Chazy XII equation after an affine change of variables:
$u^{\prime\prime\prime}+4uu^{\prime\prime}+3(u^{\prime})^{2}+6u^{2}u^{\prime}+u^{4}=0.$
Similarly, the derivative of the pure second-order Riccati equation gives a special case of the Chazy IV equation:
$u^{\prime\prime\prime}+3uu^{\prime\prime}+3(u^{\prime})^{2}+3u^{2}u^{\prime}=0.$
For $n=4$, the Riccati chain yields the Fuchs-type equation:
$u^{(4)}+5uu^{\prime\prime\prime}+10u^{\prime}u^{\prime\prime}+10u^{2}u^{\prime\prime}+15u(u^{\prime})^{2}+10u^{3}u^{\prime}+u^{5}=0,$
which coincides with the Fuchs-type equation F-XVI in the Bureau classification.
These higher-order relations for the Riccati chain and their connection to Painlevé-type equations are thoroughly explored in [12], where the stabilizer set of the Virasoro orbit and the relations between these equations are studied.
4 Indicial Loci and Kovalevskaya Exponents
We begin by recalling the autonomous part of the $m$-th Riccati chain(eq 3.2), written as an $m$-dimensional first-order system:
$\frac{dx_{i}}{dz}=f_{i}^{A}(x),\quad i=1,\dots,m,$
where the functions $f_{i}^{A}(x)$ are given by
$f_{k}^{A}(x)$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (4.1)
$f_{m}^{A}(x)$ $=-cx_{1}x_{m},$
and $c\in\mathbb{C}^{*}$ is a constant. We assign weights to the variables $x_{k}$ as follows:
$\mathrm{wt}(x_{k})=p_{k}:=k,\quad 1\leq k\leq m,$
so that the vector field $f^{A}=(f_{1}^{A},\dots,f_{m}^{A})$ is quasi-homogeneous of degree +1+1, meaning that it satisfies the scaling relation:
$f_{i}^{A}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m})=\lambda^{p_{i}+1}f_{i}^{A}(x_{1},\dots,x_{m}),\quad i=1,\dots,m,\quad\lambda\in\mathbb{C}^{*}.$
Next, we examine the formal Laurent series solutions of the truncated system of equations for small $z-z_{0}$. The solutions are of the form
$x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}+\cdots,$
where $p_{i}=i$ for $i=1,\dots,m$, and the leading coefficients $\xi=(\xi_{1},\dots,\xi_{m})$ satisfy the indicial equations
$p_{i}\xi_{i}=f_{i}^{A}(\xi),\quad i=1,\dots,m.$ (4.2)
Lemma 4.1.
For each integer $n\in\{1,\dots,m\}$, define
$\xi^{(n)}=\left(\xi^{(n)}_{1},\dots,\xi^{(n)}_{m}\right),\quad\xi^{(n)}_{k}=\begin{cases}\dfrac{(n)_{k}}{c},&1\leq k\leq n,\\ 0,&n<k\leq m,\end{cases}$
where $(n)_{k}:=n(n-1)\cdots(n-k+1)$ is the falling factorial. Then each $\xi^{(n)}$ satisfies the indicial equations (4.2). Moreover, these vectors $\xi^{(1)},\dots,\xi^{(m)}$ are the only nonzero solutions to the indicial equations.
Proof.
Substituting $x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}$ into (4.1) and comparing the leading terms, we obtain the following recursive relations:
$k\xi_{k}=\xi_{k+1}-c\xi_{1}\xi_{k},\quad 1\leq k\leq m-1,\quad-m\xi_{m}=-c\xi_{1}\xi_{m}.$
From the second equation, we have
$(c\xi_{1}-m)\xi_{m}=0.$
If $\xi_{m}\neq 0$, then $\xi_{1}=\frac{m}{c}$. Using the recursive structure of the system, we find that $\xi_{k}=\frac{(m)_{k}}{c}$ for $1\leq k\leq m$.
On the other hand, if $\xi_{m}=0$, the condition $(c\xi_{1}-(m-1))\xi_{m-1}=0$ must hold. Thus, either $\xi_{1}=\frac{m-1}{c}$ or $\xi_{m-1}=0$. If $\xi_{1}=\frac{m-1}{c}$, we obtain $\xi_{k}=\frac{(m-1)_{k}}{c}$ for $1\leq k\leq m-1$. Continuing this process, we obtain $\xi_{k}=\frac{(n)_{k}}{c}$ for some integer $n\in\{1,\dots,m\}$.
Thus, the only nonzero solutions to the indicial equations are the vectors $\xi^{(1)},\dots,\xi^{(m)}$, which correspond to the falling factorials $(n)_{k}/c$ for each $n$. ∎
Kovalevskaya matrix and principal exponents
For a given indicial locus $\xi\in{\mathbb{C}}^{m}\setminus\{0\}$, the Kovalevskaya matrix is defined as
$K(\xi):=\bigl(K_{ij}(\xi)\bigr)_{1\leq i,j\leq m},\qquad K_{ij}(\xi)=\frac{\partial f_{i}^{A}}{\partial x_{j}}(\xi)+p_{i}\delta_{ij},$ (4.3)
where $p_{i}=i$ and $\delta_{ij}$ is the Kronecker delta. Its eigenvalues are called the Kovalevskaya exponents associated with $\xi$.
Theorem 4.2.
The eigenvalues of $K(\xi^{(n)})$ are given by:
$\lambda_{k}=-n+k,\quad\text{for}\quad k=0,1,2,\dots,n-1,$
and
$\lambda_{k}=m-n+k,\quad\text{for}\quad k=n,n+1,\dots,m-1.$
Proof.
Let $K(\xi^{(n)})$ be the Kovalevskaya matrix for the indicial locus $\xi^{(n)}$. The matrix $K(\xi^{(n)})-\lambda I$ takes the form
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{m}&0&\cdots&0&m-n-\lambda\end{pmatrix}$,
where $(n)_{k}=n(n-1)\cdots(n-k+1)$ is the falling factorial. For $m>n$, the terms $(n)_{k}=0$ for $k>n$.
We begin by observing the block structure of $K(\xi^{(n)})-\lambda I$. Since $(n)_{k}=0$ for $k>n$, the matrix can be written as:
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}$,
where: - $A_{n}$ is an $n\times n$ matrix, - $C_{m-n}$ is a upper triangular matrix of size $(m-n)\times(m-n)$.
The determinant of this block matrix is given by:
$\det\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}=\det(A_{n})\cdot\det(C_{m-n}).$
The matrix $C_{m-n}$ is an upper triangular matrix with diagonal entries:
$C_{m-n}=\operatorname{diag}(m-n-\lambda,m-n+1-\lambda,\dots,m-1-\lambda).$
Thus, its determinant is:
$\det(C_{m-n})=\prod_{k=n+1}^{m}(k-n-\lambda).$
The matrix $A_{n}$ is lower triangular, and its entries are:
$A_{n}=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{n}&0&\cdots&0&-\lambda\end{pmatrix}$.
$\det(A_{n})$ $=(1-2n-\lambda)\cdot\det(A_{n}^{(1,1)})+(n)_{2}\cdot\det(A_{n}^{(1,2)})$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+\sum_{j=2}^{n-1}(n)_{j}\prod_{\begin{subarray}{c}k=j+1\end{subarray}}^{n}(k-n-\lambda)+(n)_{n}$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+(n)_{2}\prod_{\begin{subarray}{c}k=2\end{subarray}}^{n-1}(k-n-\lambda)$
$=\left((1-2n-\lambda)\lambda+(n)(n-1)\right)\prod_{k=2}^{n-1}(k-n-\lambda)$
$=\prod_{\begin{subarray}{c}k=0\end{subarray}}^{n-1}(k-n-\lambda)$
Combining the results, the determinant of $K(\xi^{(n)})-\lambda I$ becomes:
$\det(K(\xi^{(n)})-\lambda I)=\left(\prod_{k=0}^{n-1}(k-n-\lambda)\right)\cdot\left(\prod_{k=n+1}^{m}(k-n-\lambda)\right).$
To find the eigenvalues, we set the determinant equal to zero:
$\prod_{k=0}^{n-1}(k-n-\lambda)\cdot\prod_{k=n+1}^{m}(k-n-\lambda)=0.$
Thus, for each $n\in\{1,\dots,m\}$, the eigenvalues $\kappa=(\kappa_{1},\dots,\kappa_{m})$ of $K(\xi^{(n)})$ follow the pattern $(-n,-n+1,\dots,-1,1,\dots,m-n)$, which completes the proof. ∎
Definition 4.3 (Principal and lower indicial loci).
We call the indicial locus $\xi^{(1)}$ principal, since its K-exponents are exactly $\{-1,1,\dots,m-1\}$, i.e. one distinguished eigenvalue 1-1 and m1m-1 strictly positive integers. The loci $\xi^{(n)}$ with $m\geq 2$ are called lower indicial loci, in analogy with the terminology of the Painlevé test: they do not carry the full set of nonnegative integer Kovalevskaya exponents.
5 Extended Painlevé Property
The goal of this section is to prove that the Riccati chain satisfies the Extended Painlevé Property (EPP), as defined in [5]. For details on weighted projective spaces, readers can refer to [5].
The Riccati hierarchy can be understood in the context of weighted projective spaces, which are spaces where the coordinates are scaled according to specific weights under a $\mathbb{C}^{*}$-action. More specifically, the weighted projective space $\mathbb{C}P^{m+2}(1,\cdots,m,m+1,m+2)$ associated with the $m$-th Riccati chain is defined by the equivalence relation on $\mathbb{C}^{4}\setminus\{0\}$:
$(x_{1},\dots,x_{m},z,\varepsilon)\sim(\lambda x_{1},\dots,\lambda^{m}x_{m},\lambda^{m}z,\lambda^{m+1}\varepsilon),$
where $\lambda\in\mathbb{C}^{*}$. This defines the weighted projective space, where each coordinate is rescaled according to a specific weight.
The orbifold structure of this space is given by:
$\mathbb{C}P^{m+1}(1,\cdots,m,r,s)=\mathbb{C}^{m+1}/\mathbb{Z}_{r}\cup\cdots\cup\mathbb{C}^{m+1}/\mathbb{Z}_{s},$
where each $\mathbb{C}^{m+1}/\mathbb{Z}_{i}$ is a local chart for the projective space. The inhomogeneous coordinates on each chart are denoted by $(x_{1},\dots,x_{m},z)$.
The transformation between the inhomogeneous coordinates $(x_{1},\dots,x_{m},z)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{s}$ and the coordinates $(X_{1},\dots,X_{j-1},X_{j+1},\dots,X_{m},Z,\varepsilon)$ on $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ is given by:
$x_{i}=X_{i}\varepsilon^{-1/s}\quad(i\neq j),\quad x_{j}=\varepsilon^{-j/s},\quad z=Z\varepsilon^{-r/s},$
where $\varepsilon$ is a local parameter for the action of $\mathbb{C}^{*}$.
The Riccati Chain System in Local Coordinates
The Riccati chain system on the local chart $\mathbb{C}^{m+1}/\mathbb{Z}_{j}$ can be written as the following system of differential equations:
$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{d\varepsilon}=\frac{1}{s\varepsilon}\left(iX_{i}-j\frac{f_{i}+\varepsilon G_{i}}{f_{j}+\varepsilon G_{j}}\right),\quad i=1,\dots,m,\quad i\neq j,\\ \displaystyle\frac{dZ}{d\varepsilon}=\frac{1}{s\varepsilon}\left(rZ-\frac{p_{j}\varepsilon}{f_{j}+\varepsilon G_{j}}\right),\end{array}\right$. (5.1)
where $f_{i}=f_{i}(X_{1},\dots,X_{m},Z)$ are the components of the Riccati chain in the local coordinates. This system is rational and invariant under the $\mathbb{Z}_{j}$-action, reflecting the weighted projective space structure.
For convenience, we rewrite the system (5.1) as an autonomous vector field in the form:
$\left\{\begin{array}[]{l}\displaystyle\frac{dX_{i}}{dt}=iX_{i}-j\frac{X_{i+1}-cX_{1}X_{k}}{X_{j+1}-cX_{1}},\quad(i=1,\dots,m;i\neq j),\\ \displaystyle\frac{dZ}{dt}=rZ-\frac{j\varepsilon}{X_{j+1}-cX_{1}},\\ \displaystyle\frac{d\varepsilon}{dt}=s\varepsilon,\end{array}\right$. (5.2)
where the new independent variable $t$ parametrizes the integral curves of the system. This formulation makes it easier to analyze the behavior of the Riccati chain in the context of weighted projective spaces.
The following lemma describes the fixed points of the vector field in (5.2).
Lemma 5.1.
The point
$\left(X_{1},\dots,X_{m},Z,\varepsilon\right)=\begin{cases}\left(\frac{(n)_{1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-1/j},\dots,\frac{(n)_{j-1}}{c}\left(\frac{(n)_{j}}{c}\right)^{-(j-1)},0,\dots,0,0\right),&\text{for }j\leq n,\\ (0,\dots,0),&\text{for }j>n,\end{cases}$ (5.3)
is a fixed point of the vector field in (5.2).
Lemma 5.2.
The eigenvalues of the Jacobi matrix of the vector field (5.2) at the fixed point (5.3) are given by $r$, $s$, and the Kovalevskaya exponents of the system (4.1), excluding the trivial exponent 1-1. This result has been established in [5].
We now state the main result of this section:
Theorem 5.3.
The Riccati chain satisfies the Extended Painlevé Property (EPP) in the sense that it admits an mk+1m-k+1-parameter family of convergent Laurent series solutions, where the set $\{a_{i,n}\}$ includes mkm-k free parameters.
Proof.
For each indicial locus $(\xi^{(n)})$, it corresponds to n1n-1 negative eigenvalues, mnm-n positive eigenvalues, and the trivial exponent 1-1. By Lemma 4.2, the eigenvalues of the Jacobi matrix at the fixed point are given by $\kappa$, and the parameters $r$ and $s$.
By translation, the system can be decomposed into an mn+2m-n+2-dimensional unstable manifold and an n1n-1-dimensional stable manifold. We perform a suitable linear transformation, which allows us to rewrite the system in the form:
$\left\{\begin{array}[]{l}\displaystyle\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{u}\in\mathbb{C}^{m-n+2},\\ \displaystyle\frac{d\mbox{$X$}_{s}}{dt}=J_{s}\mbox{$X$}_{s}+F_{s}(\mbox{$X$}_{u},\mbox{$X$}_{s}),\quad\mbox{$X$}_{s}\in\mathbb{C}^{n-1}.\end{array}\right$.
where $J_{u}$ and $J_{s}$ are the Jacobian matrices corresponding to the unstable and stable manifolds, respectively, with the real parts of the eigenvalues of $J_{u}$ being positive and those of $J_{s}$ being negative.
By the unstable manifold theorem, there exists a local analytic function $\varphi$ such that $\varphi(0)=D\varphi(0)=0$, and the set $(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u}))$ forms the unstable manifold. The dynamics on the unstable manifold can then be expressed as:
$\frac{d\mbox{$X$}_{u}}{dt}=J_{u}\mbox{$X$}_{u}+F_{u}(\mbox{$X$}_{u},\varphi(\mbox{$X$}_{u})).$
Note that $J_{u}$ is semisimple and has distinct positive integer eigenvalues. Since for $j>n$, the fixed point is located at the origin, $(X_{1},\dots,X_{m},Z,\varepsilon)=(0,\dots,0,0)$, the fixed point lies at the origin.
Furthermore, since the system does not contain resonant terms in the Poincaré domain, we can apply Poincaré’s linearization theory to linearize the unstable manifold, completing the proof that the Riccati chain satisfies the Extended Painlevé Property. ∎
Acknowledgments. The author wishes to express sincere gratitude to Professor Hayato Chiba for his valuable guidance and instruction.