### v3

# Kovalevskaya exponents of the Riccati hierarchy

Abstract
We study the Kovalevskaya exponents of the Riccati hierarchy, deriving the general indicial loci and exponents. A unique commuting vector field is identified, and the general solution is obtained via symmetry reduction. Finally, Laurent series expansions in annular regions are analyzed to reveal the singular structure of the solutions.
1 Introduction
The study of Kovalevskaya exponents is central to understanding the integrability and singularities of nonlinear differential equations [10, 12, 15], particularly in the context of quasi-homogeneous systems [3]. These exponents offer a way to analyze the behavior of solutions to integrable systems, helping to classify and understand the singularities of their Laurent series solutions.
If we assume the existence of a hierarchy of integrable equations, it is equivalent to say that we have a family of commuting flows. In [16], we studied the commuting vector field under quasi-homogeneity settings. Motivated by this study of commuting flows in quasi-homogeneous systems and their connection to Painlevé equations, we specifically focus on the Kovalevskaya exponents of the Riccati hierarchy in this paper. The Riccati equation, a fundamental nonlinear differential equation, has higher-order generalizations known as Riccati chains. These chains naturally arise in integrability theory when taking successive covariant derivatives of a projective field, and they provide a rich structure that links linear, nonlinear, and integrable systems. The Kovalevskaya exponents of the Riccati hierarchy are crucial for determining the singular structure of these systems.
The key findings of this paper are as follows:
  • Indicial Loci and Kovalevskaya Exponents: We derive the general indicial loci and Kovalevskaya exponents for the Riccati hierarchy. These exponents are determined by a specific recursive structure, with the Kovalevskaya exponents revealing the free parameters at each stage of the hierarchy.
  • Commuting Vector Field: We identify a unique commuting vector field for the Riccati hierarchy and use symmetry reduction to derive the general solution. This reduction provides a simplified form of the solution, aiding in a clearer understanding of the structure of the solutions. In the 2-dimensional case, we also apply the blow-up method commonly used in Painlevé analysis to resolve singularities, which yields the same general solution as derived through symmetry reduction.
  • Annular Laurent Series Expansions: By analyzing Laurent series expansions in annular regions of the complex plane, we explore the singularities of the Riccati chain. We focus on how negative Kovalevskaya exponents influence the expansion, offering a deeper understanding of the structure of solutions in different regions.
This work contributes to the broader field of integrable systems by providing a detailed analysis of the Kovalevskaya exponents of the Riccati hierarchy and their connection to the singularities of the system’s solutions. Furthermore, we demonstrate how commuting vector fields and symmetry reduction techniques can be applied to simplify the solution structure, making it more tractable for further analysis. Finally, the study of Laurent series expansions in various regions of the complex plane offers new insights into the singular structure of the Riccati chain and the role of negative Kovalevskaya exponents.
2 Quasi-homogeneity of the Riccati Chain
In this section, we introduce the theoretical framework for the Riccati chain, emphasizing its recursive structure and the central concept of quasi-homogeneity. Quasi-homogeneity plays a crucial role in the integrability of differential systems, particularly in systems with the Painlevé property, by governing the scaling behavior of their solutions.
Definition 2.1.
Let $L$ be the first-order differential operator
$L=\frac{d}{dz}+cu(z),\quad c\in\mathbb{C}^{*},$
where $u(z)$ is a potential function. The $m$-order Riccati chain equation is a generalization of the standard Riccati equation and can be written as:
$L^{m}u(z)+\sum_{j=1}^{m}\alpha_{j}(z)L^{j-1}u(z)+\alpha_{0}(z)=0,$
where $m\in\mathbb{N}$, and the coefficient functions $\alpha_{j}(z)$ depend on $z$.
Next, we introduce the system of differential equations corresponding to the Riccati chain. By defining $x_{k}\rightarrow L^{k-1}u$, the Riccati chain can be written as a system of first-order differential equations:
$x_{k}^{\prime}$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (2.1)
$x_{m}^{\prime}$ $=-\sum_{j=0}^{m}\alpha_{j}(z)x_{j}-cx_{1}x_{m},\quad$ $(x_{0}\equiv 1),$
which reduces to the standard Riccati equation when $m=1$:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
We assign weights to the variables $x_{k}$ and $z$ as follows:
$\mathrm{wt}(x_{k})=p_{k}:=k,\quad 1\leq k\leq m,\quad\mathrm{wt}(z)=r\in\{1,\cdots,m+1\}$
so that the vector field $f=(f_{1},\dots,f_{m})$ is quasi-homogeneous of degree +1+1, meaning that it satisfies the scaling relation:
$f_{i}(\lambda^{p_{1}}x_{1},\dots,\lambda^{p_{m}}x_{m},\lambda^{r}z)=\lambda^{p_{i}+1}f_{i}(x_{1},\dots,x_{m},z),\quad i=1,\dots,m,\quad\lambda\in\mathbb{C}^{*}.$
The coefficient functions $\alpha_{j}(z)$ satisfy the following weight classification:
$\text{wt}(\alpha_{j})=m+1-j,\quad 0\leq j\leq m.$
In particular, if $\alpha_{j}(z)$ are polynomials in $z$, they take the form
$\alpha_{j}(z)=\begin{cases}A_{j}z^{\frac{m-j+1}{r}}&\text{if}\quad\frac{m-j+1}{r}\in\mathbb{Z},\\ 0&\text{otherwise}.\end{cases}$
Examples for $m=1$
For $m=1$, the Riccati chain reduces to the standard Riccati equation:
$x_{1}^{\prime}=-\alpha_{1}(z)x_{1}-\alpha_{0}(z)+cx_{1}^{2}.$
The weight classification of this equation gives two types of quasi-homogeneous equations:
$x_{1}^{\prime}=-z+cx_{1}^{2}\quad\text{(Airy type)},$
and
$x_{1}^{\prime}=-zx_{1}-z^{2}+cx_{1}^{2}\quad\text{(Hermite type)}.$
The Airy-type solution is discussed in [2], where a compactified Riccati equation of Airy type on a weighted projective space is studied.
Higher-order Examples and Chazy–Bureau Equations
For higher-order cases, the Riccati chain leads to more complex equations. For $m=3$, the Riccati chain in scalar form reads:
$u^{\prime\prime\prime}$ $+(\alpha_{3}(z)+4cu)u^{\prime\prime}+3c(u^{\prime})^{2}$
$+\left(6c^{2}u^{2}+3c\alpha_{3}(z)u+\alpha_{2}(z)\right)u^{\prime}$
$+c^{3}u^{4}+c^{2}\alpha_{3}(z)u^{3}+c\alpha_{2}(z)u^{2}+\alpha_{1}(z)u+\alpha_{0}(z)=0.$
For $\alpha_{j}(z)=0$ and $c=1$, this equation simplifies to the Chazy XII equation after an affine change of variables:
$u^{\prime\prime\prime}+4uu^{\prime\prime}+3(u^{\prime})^{2}+6u^{2}u^{\prime}+u^{4}=0.$
Similarly, the derivative of the pure second-order Riccati equation gives a special case of the Chazy IV equation:
$u^{\prime\prime\prime}+3uu^{\prime\prime}+3(u^{\prime})^{2}+3u^{2}u^{\prime}=0.$
For $m=4$, the Riccati chain yields the Fuchs-type equation:
$u^{(4)}+5uu^{\prime\prime\prime}+10u^{\prime}u^{\prime\prime}+10u^{2}u^{\prime\prime}+15u(u^{\prime})^{2}+10u^{3}u^{\prime}+u^{5}=0,$
which coincides with the Fuchs-type equation F-XVI in the Bureau classification.
These higher-order relations for the Riccati chain and their connection to Painlevé-type equations are thoroughly explored in [11], where the stabilizer set of the Virasoro orbit and the relations between these equations are studied.
3 Indicial Loci and Kovalevskaya Exponents
We begin by considering the autonomous part of the $m$-th Riccati chain (2.1), written as an $m$-dimensional first-order system:
$\frac{dx_{i}}{dz}=f_{i}^{A}(x),\quad i=1,\dots,m,$
where the functions $f_{i}^{A}(x)$ are given by
$f_{k}^{A}(x)$ $=x_{k+1}-cx_{1}x_{k},\quad$ $1\leq k\leq m-1,$ (3.1)
$f_{m}^{A}(x)$ $=-cx_{1}x_{m},$
and $c\in\mathbb{C}^{*}$ is a constant.
Next, we examine the formal Laurent series solutions of the truncated system of equations for small $z-z_{0}$. By quasihomogenity of the system, the solutions are of the form
$x_{i}(z)=\xi_{i}(z-z_{0})^{-p_{i}}+\cdots,$
where $p_{i}=i$ for $i=1,\dots,m$, and the leading coefficients $\xi=(\xi_{1},\dots,\xi_{m})$ are called indicial locus and satisfy the indicial equations
$p_{i}\xi_{i}=f_{i}^{A}(\xi),\quad i=1,\dots,m.$ (3.2)
Indicial loci
Lemma 3.1.
For each integer $n\in\{1,\dots,m\}$, define
$\xi^{(n)}=\left(\xi^{(n)}_{1},\dots,\xi^{(n)}_{m}\right),\quad\xi^{(n)}_{k}=\begin{cases}\dfrac{(n)_{k}}{c},&1\leq k\leq n,\\ 0,&n<k\leq m,\end{cases}$
where $(n)_{k}:=n(n-1)\cdots(n-k+1)$ is the falling factorial. Then each $\xi^{(n)}$ satisfies the indicial equations (3.2). Moreover, these vectors $\xi^{(1)},\dots,\xi^{(m)}$ are the only nonzero solutions to the indicial equations.
Proof.
Substituting $x_{i}(z)=\xi_{i}(z-z_{0})^{-i}$ into (3.2) and comparing the leading terms, we obtain the following recursive relations:
$k\xi_{k}=\xi_{k+1}-c\xi_{1}\xi_{k},\quad 1\leq k\leq m-1,\quad-m\xi_{m}=-c\xi_{1}\xi_{m}.$
From the second equation, we have
$(c\xi_{1}-m)\xi_{m}=0.$
If $\xi_{m}\neq 0$, then $\xi_{1}=\frac{m}{c}$. Using the recursive structure of the system, we find that $\xi_{k}=\frac{(m)_{k}}{c}$ for $1\leq k\leq m$.
On the other hand, if $\xi_{m}=0$, the condition $(c\xi_{1}-(m-1))\xi_{m-1}=0$ must hold. Thus, either $\xi_{1}=\frac{m-1}{c}$ or $\xi_{m-1}=0$. If $\xi_{1}=\frac{m-1}{c}$, we obtain $\xi_{k}=\frac{(m-1)_{k}}{c}$ for $1\leq k\leq m-1$. Continuing this process, we obtain $\xi_{k}=\frac{(n)_{k}}{c}$ for some integer $n\in\{1,\dots,m\}$.
Thus, the only nonzero solutions to the indicial equations are the vectors $\xi^{(1)},\dots,\xi^{(m)}$, which correspond to the falling factorials $(n)_{k}/c$ for each $n$. ∎
Kovalevskaya matrix
For a given indicial locus $\xi\in{\mathbb{C}}^{m}\setminus\{0\}$, the Kovalevskaya matrix is defined as
$K(\xi):=\bigl(K_{ij}(\xi)\bigr)_{1\leq i,j\leq m},\qquad K_{ij}(\xi)=\frac{\partial f_{i}^{A}}{\partial x_{j}}(\xi)+p_{i}\delta_{ij},$ (3.3)
where $p_{i}=i$ and $\delta_{ij}$ is the Kronecker delta. Its eigenvalues are called the Kovalevskaya exponents associated with $\xi$.
Theorem 3.2.
The eigenvalues of $K(\xi^{(n)})$ are given by:
$\lambda_{k}=-n+k,\quad\text{for}\quad k=0,1,2,\dots,n-1,$
and
$\lambda_{k}=m-n+k,\quad\text{for}\quad k=n,n+1,\dots,m-1.$
Proof.
Let $K(\xi^{(n)})$ be the Kovalevskaya matrix for the indicial locus $\xi^{(n)}$. The matrix $K(\xi^{(n)})-\lambda I$ takes the form
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{m}&0&\cdots&0&m-n-\lambda\end{pmatrix}$,
where $(n)_{k}=n(n-1)\cdots(n-k+1)$ is the falling factorial. For $m>n$, the terms $(n)_{k}=0$ for $k>n$.
We begin by observing the block structure of $K(\xi^{(n)})-\lambda I$. Since $(n)_{k}=0$ for $k>n$, the matrix can be written as:
$K(\xi^{(n)})-\lambda I=\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}$,
where: - $A_{n}$ is an $n\times n$ matrix, - $C_{m-n}$ is a upper triangular matrix of size $(m-n)\times(m-n)$.
The determinant of this block matrix is given by:
$\det\begin{pmatrix}A_{n}&0\\ 0&C_{m-n}\end{pmatrix}=\det(A_{n})\cdot\det(C_{m-n}).$
The matrix $C_{m-n}$ is an upper triangular matrix with diagonal entries:
$C_{m-n}=\operatorname{diag}(m-n-\lambda,m-n+1-\lambda,\dots,m-1-\lambda).$
Thus, its determinant is:
$\det(C_{m-n})=\prod_{k=n+1}^{m}(k-n-\lambda).$
The entries of matrix $A_{n}$ are:
$A_{n}=\begin{pmatrix}1-2n-\lambda&1&0&\cdots&0\\ -(n)_{2}&2-n-\lambda&1&\cdots&0\\ -(n)_{3}&0&3-n-\lambda&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&1\\ -(n)_{n}&0&\cdots&0&-\lambda\end{pmatrix}$.
$\det(A_{n})$ $=(1-2n-\lambda)\cdot\det(A_{n}^{(1,1)})+(n)_{2}\cdot\det(A_{n}^{(1,2)})$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+\sum_{j=2}^{n-1}(n)_{j}\prod_{\begin{subarray}{c}k=j+1\end{subarray}}^{n}(k-n-\lambda)+(n)_{n}$
$=(1-2n-\lambda)\prod_{k=2}^{n}(k-n-\lambda)+(n)_{2}\prod_{\begin{subarray}{c}k=2\end{subarray}}^{n-1}(k-n-\lambda)$
$=\left((1-2n-\lambda)\lambda+(n)(n-1)\right)\prod_{k=2}^{n-1}(k-n-\lambda)$
$=\prod_{\begin{subarray}{c}k=0\end{subarray}}^{n-1}(k-n-\lambda)$
Combining the results, the determinant of $K(\xi^{(n)})-\lambda I$ becomes:
$\det(K(\xi^{(n)})-\lambda I)=\left(\prod_{k=0}^{n-1}(k-n-\lambda)\right)\cdot\left(\prod_{k=n+1}^{m}(k-n-\lambda)\right).$
To find the eigenvalues, we set the determinant equal to zero:
$\prod_{k=0}^{n-1}(k-n-\lambda)\cdot\prod_{k=n+1}^{m}(k-n-\lambda)=0.$
Thus, for each $n\in\{1,\dots,m\}$, the eigenvalues $\kappa=(\kappa_{1},\dots,\kappa_{m})$ of $K(\xi^{(n)})$ follow the pattern $(-n,-n+1,\dots,-1,1,\dots,m-n)$, which completes the proof. ∎
Principal and lower indicial loci
Let $\kappa(\xi^{k})$ denote the Kovalevskaya exponents associated with the indicial locus $\xi^{k}$ for $k\in\{1,\dots,m\}$. We can write the Kovalevskaya exponents of Riccati chain as following:
Principal $\kappa(\xi^{1})=\{-1,1,\dots,m-1\},$
Lower $\kappa(\xi^{2})=\{-2,-1,\dots,m-2\},$
$\vdots$
Lower $\kappa(\xi^{m-1})=\{-m+1,\dots,-1,1\},$
Lower $\kappa(\xi^{m})=\{-m,\dots,-1\}.$
Definition 3.3 (Principal and lower indicial loci).
We call the indicial locus $\xi$ is principal if the associated Laurent series solution exists and includes $m$ free parameters. If the number of nonnegative integer K-exponents are smaller than m1m-1, the locus is called an lower indicial locus.
In the context of Kovalevskaya exponents of Riccati chain, the indicial locus $\xi^{(1)}$ is principal, since its K-exponents are exactly $\{-1,1,\dots,m-1\}$, i.e. one distinguished eigenvalue 1-1 and m1m-1 strictly positive integers. The loci $\xi^{(n)}$ with $n\geq 2$ are called lower indicial loci, since they do not carry the full set of nonnegative integer Kovalevskaya exponents.
We further note that in the context of Riccati chains, negative exponents are determined by the leading coefficient on each branch.
Remark 3.4.
On each Riccati-chain Laurent branch determined by $\xi=\xi^{(n)}$, the branch parameter $\xi_{1}$ determines a negative Kovalevskaya exponent n-n via $n=-c\xi_{1}$.
Corollary 3.5.
For the Riccati chain at the indicial locus $\xi^{(n)}$, the vector $\xi^{(n)}$ itself is an eigenvector of $K(\xi^{(n)})$ with eigenvalue n-n, i.e.
$K(\xi^{(n)})\,\xi^{(n)}=-n\,\xi^{(n)}.$
Proof.
In the explicit matrix form of $K(\xi^{(n)})-\lambda I$, set $\lambda=-n$. Since $(n)_{k}=0$ for $k>n$, the lower-right block does not interact with the first $n$ components. A direct substitution (using the same falling-factorial recursion as in the indicial equations) shows $(K(\xi^{(n)})+nI)\,\xi^{(n)}=0$. ∎
4 Commuting Vector Fields
Let $F=(f_{1},\dots,f_{m})$ and $G=(g_{1},\dots,g_{m})$ be quasi-homogeneous polynomial vector fields on $\mathbb{C}^{m}$. We consider the following partial differential equations:
$\frac{\partial x_{i}}{\partial z_{1}}=f_{i}(x),\quad\frac{\partial x_{i}}{\partial z_{2}}=g_{i}(x),\quad i=1,\dots,m,$
where $x=(x_{1},\dots,x_{m})\in\mathbb{C}^{m}$ and $z_{1},z_{2}\in\mathbb{C}$. We assume the following conditions:
(A1) $F$ and $G$ are quasi-homogeneous: there exist a tuple of positive integers $(a_{1},\dots,a_{m})\in\mathbb{N}^{m}$ and $\gamma\in\mathbb{N}$ such that
$f_{i}(\lambda^{a_{1}}x_{1},\dots,\lambda^{a_{m}}x_{m})$ $=\lambda^{a_{i}+1}f_{i}(x_{1},\dots,x_{m}),$
$g_{i}(\lambda^{a_{1}}x_{1},\dots,\lambda^{a_{m}}x_{m})$ $=\lambda^{a_{i}+\gamma}g_{i}(x_{1},\dots,x_{m}),$
for any $\lambda\in\mathbb{C}$. We call $\gamma$ the degree of $G$ with respect to the weight $(a_{1},\dots,a_{m})$. The degree of $F$ is assumed to be 1.
(A2) $F$ and $G$ commute with each other with respect to the Lie bracket:
$F,G]=0,\quad\text{where}\quad[F,G]_{i}=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{i}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{i}}{\partial x_{k}}(x)\right),\quad i=1,\dots,m.$
(A3) $F(x)=0$ only when $x=0$.
In the context of this paper, we consider the following proposition proposed in [16] related to the commutator and the indicial locus:
Proposition 4.1.
For a given indicial locus $c$, the identity $(K(c)+\gamma)G(c)=0$ holds. In particular, if $\gamma$ is not a K-exponent, then $G(c)=0$.
We now consider the autonomous Riccati chain $F$ on $\mathbb{C}^{m}$, where each component $f_{k}$ is given by:
$f_{k}(x)=x_{k+1}-cx_{1}x_{k},\quad 1\leq k\leq m-1,\quad f_{m}(x)=-cx_{1}x_{m},$
with $c\in\mathbb{C}^{*}$ and $x=(x_{1},\dots,x_{m})\in\mathbb{C}^{m}$. The weights of the variables are assumed to be $\mathrm{wt}(x_{k})=k$ for $k=1,\dots,m$.
Unique commuting vector field
Theorem 4.2.
Let $F$ be the autonomous Riccati chain with weights $\mathrm{wt}(x_{k})=k$. The only nontrivial quasi-homogeneous polynomial vector field commuting with $F$ is the unique commuting field $G$ of degree $\gamma=m$, given by
$g_{i}=x_{i}x_{m}\quad\text{for}\quad i<m,\quad g_{m}=x_{m}^{2}.$
For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists.
Proof.
We start by considering the quasi-homogeneous vector field $G=(g_{1},g_{2},\dots,g_{m})$, where each component $g_{i}$ is a polynomial of degree $i+\gamma$, which can be written in the general form:
$g_{i}(x)=\sum_{j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma}c_{j_{1},j_{2},\dots,j_{m}}x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{m}^{j_{m}},$
where $c_{j_{1},j_{2},\dots,j_{m}}$ are constants, and the sum is over all multi-indices $(j_{1},j_{2},\dots,j_{m})$ such that the degree constraint $j_{1}+2j_{2}+\dots+mj_{m}=i+\gamma$ is satisfied.
The commutator $F,G]_{m}(x)$ is given by:
$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{k}(x)\frac{\partial f_{m}}{\partial x_{k}}(x)\right).$
Substituting the expressions for $f_{k}(x)$ and $g_{k}(x)$, we have:
$F,G]_{m}(x)=\sum_{k=1}^{m}\left(f_{k}(x)\frac{\partial g_{m}}{\partial x_{k}}(x)-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1})\right).$
This simplifies to:
$F,G]_{m}(x)=\sum_{k=1}^{m-1}(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}(x)-cx_{1}x_{m}\frac{\partial g_{m}}{\partial x_{m}}-g_{1}(x)(-cx_{m})-g_{m}(x)(-cx_{1}).$
Now, we examine the critical term $(x_{k}-cx_{1}x_{k-1})\frac{\partial g_{m}}{\partial x_{k}}$. This term contains the factor $(x_{k}-cx_{1}x_{k-1})$, which introduces additional terms that cannot be canceled by any other terms in the expression. Specifically, when $k=m$, the structure of $g_{m}(x)=x_{m}^{2}$ creates terms that lead to an imbalance.
Thus, we find that the commutator $F,G]_{m}(x)$ cannot vanish unless $g_{i}(x)=x_{i}x_{m}$ for $i=1,\dots,m-1$ and $g_{m}(x)=x_{m}^{2}$, which is the only solution that satisfies the commutation relation.
Therefore, the only quasi-homogeneous polynomial vector fields commuting with $F$ are $F$ itself (degree 1) and the unique commuting field $G$ of degree $\gamma=m$, given by:
$g_{i}(x)=x_{i}x_{m}\quad\text{for}\quad i=1,\dots,m-1,\quad g_{m}(x)=x_{m}^{2}.$
For intermediate degrees $\gamma\in\{2,\dots,m-1\}$, no nontrivial commuting field exists. ∎
Symmetry reduction
The commuting vector field $G$ simplifies the Riccati chain by generating invariants along its flow. Specifically, along the flow of $G$, the quantities $y_{i}=\frac{x_{i}}{x_{m}}$ remain constant, which means the ratios of the $x_{i}$’s relative to $x_{m}$ are invariant. These invariants, interpreted as symmetries of the system, allow for a reduction in the dimensionality of the system. By exploiting these invariants, we reduce the Riccati chain to a simpler triangular system, where the equations are decoupled and can be solved sequentially. The general solution is then obtained in terms of rational functions of an arbitrary polynomial $u(z)$, where the solutions for $x_{1},x_{2},\dots,x_{m}$ are expressed as functions of $u(z)$. For the rest of section, we assume $c=1$ without loss of generality.
Corollary 4.3.
Let $F$ be the autonomous Riccati chain (3.1) with $c=1$, i.e.
$x_{i}^{\prime}$ $=x_{i+1}-x_{1}x_{i},\qquad$ $1\leq i\leq m-1,$ (4.1)
$x_{m}^{\prime}$ $=-x_{1}x_{m}.$
Let $G$ be the commuting vector field of degree $\gamma=m$ given in Theorem 4.2,
$g_{i}=x_{i}x_{m}\ (i<m),\qquad g_{m}=x_{m}^{2}.$
On the open set $\{x_{m}\neq 0\}$, define
$y_{i}:=\frac{x_{i}}{x_{m}}\quad(i=1,\dots,m-1),\qquad u:=\frac{1}{x_{m}}.$ (4.2)
Then the Riccati chain (4.1) is transformed into the triangular system
$y_{i}^{\prime}=y_{i+1}\quad(1\leq i\leq m-2),\qquad y_{m-1}^{\prime}=1,\qquad u^{\prime}=y_{1}.$ (4.3)
Consequently, writing $t=z-z_{0}$ and introducing constants $a_{0},\dots,a_{m-1}\in{\mathbb{C}}$, one has
$u(t)=\frac{t^{m}}{m!}+a_{m-1}\frac{t^{m-1}}{(m-1)!}+\cdots+a_{1}t+a_{0},$ (4.4)
and
$y_{i}(t)=u^{(i)}(t)\qquad(i=1,\dots,m-1).$ (4.5)
The general solution of (4.1) is therefore
$x_{m}(z)=\frac{1}{u(z-z_{0})},\qquad x_{i}(z)=\frac{y_{i}(z-z_{0})}{u(z-z_{0})}=\frac{u^{(i)}(z-z_{0})}{u(z-z_{0})}\quad(i=1,\dots,m-1).$ (4.6)
Proof.
Define $y_{i},u$ by (4.2). For $1\leq i\leq m-1$,
$y_{i}^{\prime}=\left(\frac{x_{i}}{x_{m}}\right)^{\prime}=\frac{x_{i}^{\prime}x_{m}-x_{i}x_{m}^{\prime}}{x_{m}^{2}}.$
Using (4.1), for $1\leq i\leq m-2$ we obtain
$y_{i}^{\prime}=\frac{(x_{i+1}-x_{1}x_{i})x_{m}-x_{i}(-x_{1}x_{m})}{x_{m}^{2}}=\frac{x_{i+1}}{x_{m}}=y_{i+1},$
and for i=m1i=m-1,
$y_{m-1}^{\prime}=\frac{x_{m}}{x_{m}}=1.$
Next,
$u^{\prime}=\left(\frac{1}{x_{m}}\right)^{\prime}=-\frac{x_{m}^{\prime}}{x_{m}^{2}}=-\frac{-x_{1}x_{m}}{x_{m}^{2}}=\frac{x_{1}}{x_{m}}=y_{1}.$
This proves (4.3).
Integrating (4.3) yields $y_{m-1}(t)=t+a_{m-1}$, then $y_{m-2}(t)=\frac{t^{2}}{2!}+a_{m-1}t+a_{m-2}$, and inductively $y_{1}$ is a polynomial of degree m1m-1 whose coefficients are affine functions of the constants $a_{0},\dots,a_{m-1}$.
$y_{m-k}(t)=\frac{t^{k}}{k!}+\frac{a_{m-1}t^{k-1}}{(k-1)!}+\frac{a_{m-2}t^{k-2}}{(k-2)!}+\cdots+a_{m-k},\quad k=1,\dots,m-1.$
Finally, $u^{\prime}=y_{1}$ implies that $u$ is a polynomial of degree $m$, given by
$u(t)=\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}.$
The identities $y_{i}=u^{(i)}$ follow by differentiating $u^{\prime}=y_{1}$ repeatedly, hence (4.6) follows from $x_{m}=1/u$ and $x_{i}=y_{i}/u$. ∎
In particulae, the general solution $x_{1}(z)$ is written as
$x_{1}(z)$ $=\frac{y_{1}(z-z_{0})}{u(z-z_{0})}$ (4.7)
$=\frac{u^{(1)}(z-z_{0})}{u(z-z_{0})}$
$=\frac{\frac{t^{m-1}}{m-1!}+\frac{a_{m-1}t^{m-2}}{(m-2)!}+\cdots+a_{1}}{\frac{t^{m}}{m!}+\frac{a_{m-1}t^{m-1}}{(m-1)!}+\cdots+\frac{a_{1}t}{1!}+a_{0}}$
$=\frac{mt^{m-1}+m(m-1)a_{m-1}t^{m-2}+\cdots+m!a_{1}}{t^{m}+ma_{m-1}t^{m-1}+\cdots+m!a_{0}}$
$=\frac{mt^{m-1}+(m-1)b_{m-1}t^{m-2}+\cdots+(m-1)!b_{1}}{t^{m}+b_{m-1}t^{m-1}+\cdots+b_{0}}$
where $b_{m-k}=m(m-1)\cdots(m-k+1)a_{m-k}$.
Blow-up method
In the context of the 2-dimensional Riccati chain, we apply the blow-up method commonly used in Painlevé analysis to resolve singularities[13, 14]. This procedure provides a regular foliation on a smooth model and will be shown to yield the same general solution as derived through symmetry reduction. The blow-up method introduces a sequence of transformations that compactify the space and resolve indeterminacies at infinity, leading to a simpler structure that can be analyzed more easily.
Compactification and a base point at infinity
Consider the $2$-dimensional autonomous system
$x_{1}^{\prime}=x_{2}-x_{1}^{2},\qquad x_{2}^{\prime}=-x_{1}x_{2}.$ (4.8)
On $\mathbb{P}$$\times{\mathbb{P}}$ introduce the chart at infinity
$(X,Y):=\Bigl(\frac{1}{x_{1}},\frac{1}{x_{2}}\Bigr),$
so that the induced vector field is
$X^{\prime}=1-\frac{X^{2}}{Y},\qquad Y^{\prime}=\frac{Y}{X}.$
The point $(X,Y)=(0,0)$ is a base point, since the right-hand sides are indeterminate there.
Blow-up
Blow up $(X,Y)=(0,0)$. In the chart X=u,Y=uvX=u,\ Y=uv one obtains
$u^{\prime}=\frac{v-u}{v},\qquad v^{\prime}=1.$ (4.9)
Equation (4.9) is integrable:
$v(z)=z-z_{0},\qquad u(z)=\frac{(z-z_{0})^{2}+2C}{2(z-z_{0})}=\frac{z^{2}-2zz_{0}+z_{0}^{2}+2C}{2z-2z_{0}},$
where $C$ is an arbitrary constant.
With $b_{1}=2z_{0}$ and $b_{2}=z_{0}^{2}+2C$, we recover the same general solution as (4.7) after transforming back to $(x_{1},x_{2})$.
5 Annular Laurent series expansions
In this section, we examine the Laurent expansion of the solution $x_{1}(z)$ in the annular regions of the complex plane, and explore how the negative Kovalevskaya exponent n-n varies with the choice of region (disk, annulus, or exterior) for the expansion. We begin by rewritting our general solution (4.7) into a sum over simple poles:
$x_{1}(z)=\sum_{j=1}^{m}\frac{1}{z-z_{j}},$ (5.1)
where $z_{1},\dots,z_{m}$ are the singularities of the solution.
Next, we introduce the transformation $w:=z-z_{1}$ and $\delta_{j}:=z_{j}-z_{1}$ for $j\geq 2$, and rewrite the solution as:
$x_{1}(w)=\frac{1}{w}+\sum_{j=2}^{m}\frac{1}{w-\delta_{j}}.$ (5.2)
For each term $\frac{1}{w-\delta_{j}}$, we examine the geometric series expansions depending on the region where the expansion is performed.
Geometric series expansion
We consider the following two cases based on the relative magnitudes of $|w|$ and $|\delta_{j}|$:
  • Inside the disc ($|w|<|\delta_{j}|$): In this case, the geometric series expansion is:
    $\frac{1}{w-\delta_{j}}=-\frac{1}{\delta_{j}}\cdot\frac{1}{1-w/\delta_{j}}=-\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}},$ (5.3)
    which does not contain a $w^{-1}$ term.
  • Outside the disc ($|w|>|\delta_{j}|$): Here, the expansion becomes:
    $\frac{1}{w-\delta_{j}}=\frac{1}{w}\cdot\frac{1}{1-\delta_{j}/w}=\frac{1}{w}\sum_{k=0}^{\infty}\Big(\frac{\delta_{j}}{w}\Big)^{k},$ (5.4)
    where the leading term is $1/w$.
Annular Regions
Let $0<r_{2}\leq r_{3}\leq\cdots\leq r_{m}$ be the ordered moduli of $\delta_{j}$ (i.e. the radii $|\delta_{j}|$ in nondecreasing order). Define annular regions
$A_{n}:=\{w\in{\mathbb{C}}:\ r_{n}<|w|<r_{n+1}\},\qquad n=1,\dots,m,$ (5.5)
with the conventions $r_{1}:=0$ and $r_{m+1}:=+\infty$. Thus: $A_{1}$ is a punctured disc around $w=0$, intermediate $A_{n}$ are annuli, and $A_{m}$ is the exterior region.
For a fixed annulus $A_{n}$, split the set of poles other than $z_{1}$ into
$I_{\rm in}(n):=\{\,j\in\{2,\dots,m\}:\ |\delta_{j}|<|w|\ \text{for }w\in A_{n}\,\},\qquad I_{\rm out}(n):=\{2,\dots,m\}\setminus I_{\rm in}(n).$
By construction, $|I_{\rm in}(n)|=n-1$ and $|I_{\rm out}(n)|=m-n$.
Proposition 5.1.
Fix $n\in\{1,\dots,m\}$ and an annulus $A_{n}$ as in (5.5). Then $x_{1}(w)$ admits on $A_{n}$ a Laurent expansion of the form
$x_{1}(w)=\frac{n}{w}+H_{n}(w),$ (5.6)
where $H_{n}(w)$ is holomorphic on $A_{n}$ and admits a Laurent series that converges normally on $A_{n}$. Specifically, the explicit form of $H_{n}(w)$ is:
$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}\;-\;\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$ (5.7)
In particular, the coefficient of $w^{-1}$ in this Laurent expansion is $n$.
Proof.
We begin by recalling the expression for $x_{1}(w)$ given in (5.2). For each $j\in I_{\rm in}(n)$, we have $|\delta_{j}/w|<1$, so the series expansion (5.3) converges absolutely and uniformly on compact subsets of $A_{n}$. Similarly, for each $j\in I_{\rm out}(n)$, we have $|w/\delta_{j}|<1$, so the series (5.4) converges absolutely and uniformly on compact subsets of $A_{n}$.
Substituting these expansions into (5.2) yields:
$x_{1}(w)=\frac{1}{w}+\sum_{j\in I_{\rm in}(n)}\left(\frac{1}{w}+\sum_{k\geq 1}\frac{\delta_{j}^{k}}{w^{k+1}}\right)+\sum_{j\in I_{\rm out}(n)}\left(-\sum_{k\geq 0}\frac{w^{k}}{\delta_{j}^{k+1}}\right).$
Thus, the coefficient of $w^{-1}$ is computed as:
$1+|I_{\rm in}(n)|=1+(n-1)=n.$
The remaining terms form the Laurent series $H_{n}(w)$, which is holomorphic on $A_{n}$ and has the explicit form:
$H_{n}(w)=\sum_{j\in I_{\rm in}(n)}\sum_{k=1}^{\infty}\frac{\delta_{j}^{k}}{w^{k+1}}-\sum_{j\in I_{\rm out}(n)}\sum_{k=0}^{\infty}\frac{w^{k}}{\delta_{j}^{k+1}}.$
Remark 5.2.
By Proposition 5.1, the negative Kovalevskaya exponent can be interpreted as the annulus index (disk/annulus/exterior), which indicates the residue of $x_{1}$ at the movable pole. For $m\geq 3$, there are $m$ regions $A_{n}$, and thus multiple admissible Laurent expansions around a given movable pole. The inner expansion corresponds to $A_{1}$ (the punctured disk), the outer expansion corresponds to $A_{m}$ (the exterior region), and the intermediate annuli $A_{n}$ provide Laurent expansions with contributions from both inner and outer regions, with the leading coefficient being $n\in\{1,\dots,m\}$.
Acknowledgments. The author wishes to express sincere gratitude to Professor Hayato Chiba for his valuable guidance and instruction.